OpenStudy (bluesea24):

why is value of sin 45 degree always equal to 1/sq rt(2)

OpenStudy (anonymous):

http://www.intmath.com/blog/how-do-you-find-exact-values-for-the-sine-of-all-angles/6212 visit this website and you will get your answer

OpenStudy (zehanz):

You can leave out the "always". It's the same for every other angle: every angle comes with a sine value that is ...always... the value it is. E.g. sin 30 deg = 1/2, last year, now and for many years to come ;) Otherwise, use @HappySoul's hint!

OpenStudy (mathstudent55):

 The coordinates of point (x, y) must satisfy the equation x^2 + y^2 = 1 since point (x, y) lies on the unit circle. Since the anlge is 45 degrees, the triangle is an isoscleles triangle, and x = y, so you end up with y^2 + y^2 = 1 2y^2 = 1 y^2 = 1/2 y = 1/sqrt(2) The sin45 = y/1 = 1/sqrt(2)

OpenStudy (bluesea24):

hey i think now i got it....thanks a lot for those ans :)